OpenStudy (anonymous):

Find the exact time it takes a stone to reach the ground if it is dropped from a bridge that is 200 feet high.

OpenStudy (lgbasallote):

have you tried using t = \(\large \sqrt{\frac{2d}{a}}\)

OpenStudy (anonymous):

First what you need to know is that the stone will roughly fall at 32 ft/sec. After you know that you can use one of the kinematic equations to find the answer you're looking for. d=1/2at^2. So, 200=1/2(32)(t^2) Solve for t. And you get 3.53 seconds roughly

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

@marvin.cling can you write this in a simplified form? t=\[\sqrt{h/16}\]

OpenStudy (anonymous):

t=\[\sqrt{400/32}\ or 5\[\sqrt{2}\]2

OpenStudy (anonymous):

*\[5\sqrt{2}/2\]